The lengths of the legs of an isosceles triangle are integers. The base is half as long as each leg. What are the possible lengths of the legs if the perimeter is between 6 units and 16 units?
Answer by stanbon(75887)

The lengths of the legs of an isosceles triangle are integers. The base is half as long as each leg. What are the possible lengths of the legs if the perimeter is between 6 units and 16 units?
------
Let the equal sides be x and x.
The base is x/2
----
Equation:
x + x + x/2 = 16
----
(5/2)x = 16
x = (2/5)(16)
--
x = 32/5 or 6 2/5 units
----
x/2 = 16/5 or 3 1/5 units
----------
Cheers,
Stan H.
------------